OpenStudy (anonymous):

Tiffany kicks a soccer ball off the ground and in the air, with an initial velocity of 28 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 15.6 feet 14.1 feet 13.5 feet 12.3 feet

OpenStudy (igreen):

\(\sf H(t) = -16t^2 + vt + s\) Okay, plu gin what we know: \(\sf H(t) = -16t^2 + 28t + 0\) Now find the vertex, which will be the maximum height.

OpenStudy (anonymous):

Yet again IGreen beats me to the answer :P

OpenStudy (igreen):

Formula to find the x-value: \(\sf \dfrac{-b}{2a}\)

OpenStudy (igreen):

Our equation is in the form of: \(\sf ax^2 + bx + c\) So in this case: \(\sf a = -16\) \(\sf b = 28\)

OpenStudy (igreen):

@GhostVortex :P

OpenStudy (anonymous):

yea um so do you add 16 to the other side and then subtract the 28

OpenStudy (igreen):

So we have: \(\sf \dfrac{-b}{2a}\) \(\sf \dfrac{-28}{2(-16)}\) Simplify

OpenStudy (anonymous):

so you dont

OpenStudy (igreen):

No..just multiply 2 * -16 and divide.

OpenStudy (igreen):

Then we plug it back into the equation to find the maximum height.

OpenStudy (igreen):

Tell me what you get.

OpenStudy (chosenmatt):

yup

OpenStudy (anonymous):

28/-32

OpenStudy (anonymous):

7/8

OpenStudy (anonymous):

but i dont think you cn really divide that

OpenStudy (igreen):

Yes, you can.. 7 / 8 = 0.875

OpenStudy (igreen):

Just plug it in your calculator..anyway that's 't', so we plug it back into our equation.

OpenStudy (igreen):

\(\sf H(t) = -16t^2 + 28t\) \(\sf H(t) = -16(0.875)^2 + 28(0.875)\) Can you simplify that?

OpenStudy (anonymous):

yes its 12.25

OpenStudy (igreen):

Yep! That's our answer.

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

@iGreen can u help me with another one

OpenStudy (igreen):

I might be able to.

OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 2x = 17, rounded to the nearest hundredth? x ≈ -2.33 and x ≈ 1.83 x ≈ -1.83 and x ≈ 2.33 x ≈ -9.31 and x ≈ 7.31 No real solutions

OpenStudy (anonymous):

@kate_liller

OpenStudy (anonymous):

4x² + 2x - 17 = 0 x = -2/8 ± √(4 + 16•17) / 8 x = -1/4 ± √(1 + 4•17) / 4 x = -0.25 ± 2.0767 x ≈ 1.83 or x ≈ -2.33 which is the 1st choice

OpenStudy (anonymous):

thxs so much kate

OpenStudy (anonymous):

@haydentroy can u give her a medal for me